# QSO MUSEUM II: Search for extended Ly emission around 8 quasar pairs

### v1

@@ 4.1 Surface brightness levels of extended Ly emission
As described in the previous section, the PSF and continuum subtracted datacubes should only contain nebular emission lines and noise. In this work we only focus on the extended Ly emission surrounding the quasar pairs. To facilitate comparisons with previous studies, for the subsequent analysis we choose pseudo narrowbands (PNBs) with a width of 30 in observed frame around the Ly nebula line. To select the central wavelength of the PNBs, we proceed as follows. After constructing a PNB centered on the peak of the nebula emission line determined by visual inspection, we fit a Gaussian function to the spectrum extracted within the largest 2 isophote close to the respective quasar, determined from the surface brightness limit (SBL) in this PNB. We then construct a new PNB centered on the peak of the fitted line and iterate the process. We select those that maximize the flux of the fitted line as optimized PNBs. Table 2 summarizes the properties of such optimized PNBs for each pair, and Appendix A provides an example of a PNB before PSF subtraction.
@@
To account for narrower or wider lines and different nebula redshifts in cubes with two quasars in the FoV, we performed the same iteration individually for each quasar while also changing the width of the PNB to two times the FWHM of the Gaussian fit. This does not significantly change the resulting PNB or sensitivity, as the narrowbands optimized in this way are mostly already 30 wide, encompassing both nebulae. The only exception is quasar pair 2 (Fig. 4), where the determined peak wavelengths are offset from each other by 7.5 . The obtained PNBs for all quasar pairs are shown as surface brightness maps in the first panel of Figs. 3 to 10.
@@
Additionally, we compute the mean surface brightness in logarithmically spaced annuli around the brighter quasar in each pair, extending out to 450 comoving kpc or 110 physical kpc at the typical redshift of our sample. During this profile calculation, residual background emission is subtracted by calculating the mean surface brightness within a sigma-clipped background region of each image. This calculation is performed both for the full PNB and for the PNB after masking regions outside the 2 isophote shown in Figs. 3 to 10. These latter profiles are typically brighter as they do not include regions with no signal in the calculation.
@@
To get the typical surface brightness profile of physical quasar pairs, we stack the individual profiles, putting the brighter quasar in the pair to 0 ckpc and taking the median value for each radius (Fig. 11, Table 3). We made this choice because we always find brighter emission closer to the brighter quasar of the pair even though in quasar pair 3 the two AGNs have similar magnitudes and one does not show associated extended emission. For comparison, the median surface brightness profiles of other quasar samples are shown. These can typically be described with a power law with (e.g., Borisova et al. 2016; Arrigoni Battaia et al. 2019a). However, in the case of quasar pairs, the surface brightness starts out with relatively low values close to the brighter quasar and falls off more gradually. A fitted power law to the profile weighted by the symmetrized 25th and 75th percentiles yields . Fitting is performed using the nonlinear least squares function curve_fit in scipy. As the slope might be influenced by the presence of dense gas in the host halo of the companion quasar, we repeat this analysis by only stacking the three pairs at wide separation. The obtained profile is not contaminated by a second quasar halo and is best fitted with a power law index of for the median stacked profile and for the mean profile.
@@ 4.2 Ly nebulae spectra and kinematics
We obtained a global nebula spectrum for each detection by adding the emission from spaxels within the largest 2 isophote in the psf-subtracted surface brightness maps close to the respective quasars. These spectra, together with the corresponding quasar spectra extracted within a aperture centered on each quasar, are displayed in the fourth panel of Figs. 3 to 10. As the area over which the nebula spectra are integrated can be quite large and the targeted quasars are relatively faint, the peak of the nebular emission line is in some cases brighter than the peak of one of the quasar’s emission at the nebula peak wavelength (i.e., quasar pair 1, 3, 4, and 8).
From these Ly nebula spectra we can obtain a first probe of the cool gas kinematics, even though such integrated spectra encompass large portions of the quasars’ surroundings, mixing kinematics on different scales. We calculate the second order of the nebular emission line, , in a window of 2 FWHM obtained from a Gaussian fit to each spectrum. Errors are obtained by resampling the spectra with the associated noise spectra 10000 times and calculating the second moment. The 32nd and 68th percentiles of the obtained distribution are assumed as lower and upper error. The mean value of of all nebulae, 434 km s-1, is indicative of a relatively quiescent cool CGM gas, as opposed to violent kinematics with typical linewidths and velocity shifts above 1000 km s-1 (Villar-Martín et al., 2003).
To further analyze the Ly kinematics in each nebula, we obtain the line velocity shift and velocity dispersion by calculating the first and second moment from the subcubes (i.e., wavelength range) used to build the PNBs. In particular, for each pair, we set the central slice of the PNB wavelength range as reference velocity of the nebula and compute the moments within the 2 isophote. The obtained maps are shown in the second and third panels of Figs. 3 to 10. We find that the detected nebulae have velocity shifts of a few hundred km s-1, and average velocity dispersions of 390 km s-1.
@@ 4.3 Ly nebulae morphology
To quantify the morphology of the detected nebulae, we determine multiple metrics using the largest 2 isophote associated with the respective quasar: the integrated Ly nebula luminosity , the maximum extent of the nebula as well as the maximum extent between quasar position and nebula 2 isophote, (visualized in Fig. 6), the offset between quasar and nebula centroid , the enclosed area , the ratio between major and minor axis , and the angular offset between the nebula semimajor axis and the vector connecting the two quasars in the pair, . To determine and , as described in Arrigoni Battaia et al. (2019a), we first calculate the second order moments of the flux distribution as
@@
with the fluxweighted centroid coordinates and the distance of the point to the centroid, . From this, the Stokes parameters follow as
and are used to calculate the axis ratio, or asymmetry of the nebula,
and the angle of the semimajor axis to the next x- or y-axis,
@@
From , the angle between semimajor axis of the nebula and the connecting line between the quasars can be determined. The determination of positions and as a result most aforementioned values are seeing-limited and this imprecision accounts for the majority of uncertainties. We therefore assume an error of 1seeing on all distance measurements and obtain errors for and by randomly varying the centroid position 500 times within a circle of radius 1seeing. The minimum and maximum values of the obtained distribution are taken as upper and lower errors. Physical area measurements are also influenced by uncertainties on the redshift. A conservative estimate for this error is 10 percent of the measured area (Arrigoni Battaia et al., 2023b). To avoid taking into account the same nebula twice, we only determine morphologies with respect to the brighter quasar in the pair if both of them are associated with the same nebula or only one quasar is in the FoV (ID 1.1, 4.1, 5.1, 6.1 and 8.1). If there are two clearly distinct nebulae or only one nebula, analysis is performed for each of those individually with respect to the quasar they are associated with (ID 2.1, 2.2, 3.2, 7.1). This analysis is not performed for quasar 7.2 as we require a minimum enclosed area of 2 arcsec2 to be able to sensibly constrain the morphology. Table 4 lists all the aforementioned metrics computed to constrain the morphology of the detected Ly nebulae. On average, the extended emission spans about 90 kpc with a nebula luminosity of 2.8 erg s-1.
@@
We summarize our main findings on nebulae morphologies in Figs. 12 and 13. The first figure shows the axis ratio of each nebula, , against its nebula luminosity , with the dashed red line indicating the median value of . For comparison, we also provide the values reported in Arrigoni Battaia et al. (2019a) for nebulae detected around more luminous single quasars (QSO MUSEUM I) together with their median value (turquoise dash-dotted) and their 25th and 75th percentiles, and the median value of found for nebulae associated with dimmer quasars (dotted gray line, Mackenzie et al. 2021). For a better visualisation of the data, the side plots of Fig. 12 provide the distributions of all the points as histograms.
@@
We find that the second least luminous nebula considered for the morphological calculations, ID 2.1, has the most circular shape. The rest of the sample can be divided into two groups: The first one falls well within the distribution of QSO MUSEUM I with moderate asymmetries (), while the second group shows exceptionally high asymmetries (). Spanning almost two orders of magnitude in nebula luminosities and encompassing quasars 2.2, 7.1 and 8, and therefore a multitude of projected separations and magnitudes; there is no clear trend in what drives these asymmetries. Overall, the shape of quasar pair nebulae is more asymmetric than single quasar nebulae with a median value below the 25th percentile of bright quasar nebulae, even though in previous studies, there is a trend of increasing with lower quasar luminosity. In other words, we find that quasars with luminosities as low as those in Mackenzie et al. (2021), but being part of a pair, are associated with more asymmetric nebulae. The integrated nebula luminosity spans a similar range in both samples, but peaks at a higher value for single quasars, with the brightest nebula having a luminosity of 1.8 erg s-1, and a higher average luminosity of 4.5 erg s-1. Such more asymmetric morphologies in quasar pairs could be due to the presence of intergalactic structures connecting the two quasars in each pair. This tentative evidence of more asymmetric morphologies needs to be confirmed by a larger sample of quasar pair observations.
@@
To investigate this further, Fig. 13 displays the relation between (the maximum distance between the bright quasar in each pair and its associated nebula isophote) and (the angle between the nebula semimajor axis and the line connecting the two quasars). In the case of isolated/unrelated quasars we expect to find a random distribution of angles and no trend with distance. Instead, the two values are correlated with a Spearman correlation coefficient and a chance of coincidental correlation of . We further test the null hypothesis by randomly orienting the nebulae 50000 times and calculating the Spearman coefficient for the obtained distributions. In 0.7 % of realisations, the Spearman coefficient is at least as significant as for the data distribution, that is, below -0.7833, and in 0.2 % of realisations, the 4 most extended nebulae all show alignment angles below 20 degrees.
@@
The correlation is not driven by the magnitude of the quasars (color of the dots) or the projected distance between the pair (size of the dots), although at lower projected distances, the offset angle tends to be smaller. Possibly due to the small sample size, the significance of this trend is low, with a -value of 0.5 determined by performing a 2D Kolmogorov–Smirnov test (Peacock, 1983; Fasano & Franceschini, 1987; Press et al., 2007) using the public code ndtest444Written by Zhaozhou Li, https://github.com/syrte/ndtest.
@@ 4.4 Notes on each quasar pair
Quasar 1.1 (Fig. 3) is the brightest object of the sample, being as bright as the quasars targeted in Borisova et al. (2016) or Arrigoni Battaia et al. (2019a). It is also associated with the brightest Ly nebula with a luminosity . Accordingly, its Ly surface brightness profile is an order of magnitude brighter than the median profile of all quasar pairs. Quasar 1.2, on the other hand, is the faintest targeted object and sits at a projected distance of 47.7 kpc from quasar 1.1. The Ly nebula shows a slight elongation between the two quasars, also indicated by the low value of , and the nebular line velocity shift evolves from around quasar 1.1 to close to quasar 1.2. However, this is not reflected in the systemic redshifts of the AGNs (see top-middle panel in Fig. 3), likely due to the associated large uncertainties on these values as only broad lines are available for the redshift determination, leading to typical errors of up to (Shen et al., 2016). The extended Ly emission traces rather quiescent kinematics with an average velocity dispersion of 385 km s-1.
@@
Quasar pair 2, displayed in Fig. 4, is associated with the smallest Ly nebulae in both area and extent. The extended Ly emission around quasar 2.1 is also the only one in the sample that is almost circularly shaped according to the asymmetry parameter and therefore, the value of , evaluated here to be about 40∘, is not well-constrained. The two nebulae do not seem to be physically connected, supported by their redshift, which is close to the systemic redshift of the respective quasar. This might indicate that, although the projected separation is relatively small (100 kpc), the quasars could be at a significant distance from each other. In this framework the pair would be stretched along our line of sight out to a distance of 2.7 Mpc obtained by assuming that the velocity difference between the two quasars ( km s-1) is all due to the Hubble flow.
Alternatively, the two quasars might not be
able to illuminate gas in the transverse direction due to obscuration. In other words, their ionization cones do not illuminate the gas in between the two objects. While quasar 2.1 and 2.2. are relatively faint, there have been observations of individual quasars of similar luminosities which show both similar or more extended associated nebulae (e.g., compare ID4, ID5, ID6 and ID7 in Mackenzie et al. 2021).
@@
Quasar 3.1 and 3.2 have similar absolute magnitudes and are the second faintest objects (after ID 1.2) observed in this study, but are associated with the third-biggest nebula, which is displayed in Fig. 5. We note that this is true even when taking into account the slightly differing PSF subtraction method used for this source which impacts the estimated . Indeed, the standard subtraction of an empirical PSF leads to artifacts close to the quasar, reducing the area enclosed by the 2 isophote, but does not influence different measures of nebula extent like , which are sensitive to large scale emission. The discovered extended emission does not stretch between the quasars, akin to other pairs in the sample, but is concentrated around quasar 3.2 and shows relatively quiescent kinematics ( km s-1) and a shallower surface brightness profile with respect to the median profile of the sample. While the two quasars are similarly luminous, in the spectrum of quasar 3.2, a blueshifted absorption trough is visible in the CIV line (Fig. 1), which may indicate a nuclear outflow (Weymann et al. 1991) possibly helping in illuminating the surrounding gas (Costa et al. 2022).
@@
Displayed in Fig. 7 are the results for quasar 5.1, while quasar 5.2 has not been observed due to the high angular separation. Quasar 5.1 is closely accompanied by two continuum sources first detected in the MUSE data cube and examined in more detail in a companion paper (Herwig et al. in prep.). The detected extended emission is found at a significant offset from the quasar, , and is not directly connected to it. The velocity shift in this nebula spans a large range compared to the rest of the sample, with an evolution from to in East-West direction. The cool gas traced by the Ly emission might be expelled from the interstellar medium of the quasar and companion host galaxies due to tidal interactions.
@@
This object, shown in Figue 10, has already been published in Arrigoni Battaia et al. (2019b) and is reanalyzed in this work to ensure consistency. While the overall shape of the extended emission stays constant in both works, the method employed here is more conservative (pseudo-NB vs channel-by-channel detection) and consequently does not detect low surface brightness emission features. While the nebulae associated with the two quasars are still connected through a gaseous filament, the emission is much less coherent and smaller in area. However, as the emission appears physically connected, we treat it as one nebula in the analysis.

### v2

@@ 4.1 Surface brightness levels of extended Ly emission
As described in the previous section, the PSF and continuum subtracted data cubes should only contain nebular emission lines and noise. In this work we only focus on the extended Ly emission surrounding the quasar pairs. To facilitate comparisons with previous studies, for the subsequent analysis we chose PNBs with a width of 30 in observed frame around the Ly nebula line. To select the central wavelength of the PNBs, we proceeded as follows. After constructing a PNB centered on the peak of the nebula emission line determined by visual inspection, we fitted a Gaussian function to the spectrum extracted within the largest 2 isophote close to the respective quasar, determined from the SBL in this PNB. We then constructed a new PNB centered on the peak of the fitted line and iterated the process. We selected those that maximize the flux of the fitted line as optimized PNBs. Table 1 summarizes the properties of such optimized PNBs for each pair, and Appendix B provides an example of a PNB before PSF subtraction as well as the corresponding example of SB profile extracted from the PNB before PSF subtraction, after PSF subtraction, and of the empirical PSF.
@@
To account for narrower or wider lines and different nebula redshifts in cubes with two quasars in the FoV, we performed the same iteration individually for each quasar while also changing the width of the PNB to two times the FWHM of the Gaussian fit. This does not significantly change the resulting PNB or sensitivity, as the narrowbands optimized in this way are mostly already 30 wide, encompassing both nebulae. The only exception is quasar pair 2 (Fig. 7), where the determined peak wavelengths are offset from each other by 7.5 . The obtained PNBs for all quasar pairs are shown as surface brightness maps in the first panel of Figs. 3 and 7 to 13.
@@
Additionally, we computed the mean surface brightness in logarithmically spaced annuli around the brighter quasar in each pair, extending out to 450 comoving kiloparsecs or 110 physical kiloparsecs at the typical redshift of our sample. During this profile calculation, residual background emission was subtracted by calculating the mean surface brightness within a sigma-clipped background region of each image. This calculation was performed both for the full PNB and for the PNB after masking regions outside the 2 isophote shown in Figs. 3 and 7 to 13. The second set of profiles are typically brighter, as they do not include regions with no signal in the calculation.
@@
To get the typical surface brightness profile of physical quasar pairs, we stacked the individual profiles, putting the brighter quasar in the pair to 0 ckpc and taking the median value for each radius (Fig. 4, Table 2). We made this choice because we always find brighter emission closer to the brighter quasar of the pair even though in quasar pair 3 the two AGNs have similar magnitudes and one does not show associated extended emission. For comparison, the median surface brightness profiles of other quasar samples are shown. These can typically be described with a power law with (e.g., Borisova et al. 2016; Arrigoni Battaia et al. 2019a). However, in the case of quasar pairs, the surface brightness starts out with relatively low values close to the brighter quasar and falls off more gradually. A fitted power law to the profile weighted by the symmetrized 25th and 75th percentiles yields . We performed fitting using the nonlinear least squares function curve_fit in scipy (Virtanen et al., 2020). As the slope might be influenced by the presence of dense gas in the host halo of the companion quasar, we repeated this analysis by only stacking the three pairs at wide separation. The obtained profile is not contaminated by a second quasar halo and is best fitted with a power law index of for the median stacked profile and for the mean profile.
@@ 4.2 Ly nebulae spectra and kinematics
We obtained a global nebula spectrum for each detection by adding the emission from spaxels within the largest 2 isophote in the PSF-subtracted surface brightness maps close to the respective quasars. These spectra, together with the corresponding quasar spectra extracted within a aperture centered on each quasar, are displayed in the fourth panel of Figs. 3 and 7 to 13. As the area over which the nebula spectra are integrated can be quite large and the targeted quasars are relatively faint, the peak of the nebular emission line is in some cases brighter than the peak of one of the quasar’s emission at the nebula peak wavelength (i.e., quasar pair 1, 3, 4, and 8).
From these Ly nebula spectra we can obtain a first probe of the cool gas kinematics, even though such integrated spectra encompass large portions of the quasars’ surroundings, mixing kinematics on different scales. We calculated the second order of the nebular emission line, , in a window of 2 FWHM obtained from a Gaussian fit to each spectrum. Errors were obtained by resampling the spectra with the associated noise spectra 10,000 times and calculating the second moment. The 32nd and 68th percentiles of the obtained distribution were assumed as lower and upper error. The mean value of of all nebulae, 434 km s-1, is indicative of a relatively quiescent cool CGM gas, as opposed to violent kinematics with typical linewidths and velocity shifts above 1000 km s-1 (Villar-Martín et al., 2003).
To further analyze the Ly kinematics in each nebula, we obtained the line velocity shift and velocity dispersion by calculating the first and second moment from the subcubes (i.e., wavelength range) used to build the PNBs. In particular, for each pair, we set the central slice of the PNB wavelength range as reference velocity of the nebula and computed the moments within the 2 isophote. The obtained maps are shown in the second and third panels of Figs. 3 and 7 to 13. We find that the detected nebulae have velocity shifts of a few hundred km s-1, and average velocity dispersions of 390 km s-1.
@@ 4.3 Ly nebulae morphology
To quantify the morphology of the detected nebulae, we determined multiple metrics using the largest 2 isophote associated with the respective quasar: the integrated Ly nebula luminosity , the maximum extent of the nebula as well as the maximum extent between quasar position and nebula 2 isophote, (visualized in Fig. 9), the offset between quasar and nebula centroid , the enclosed area , the ratio between major and minor axis , and the angular offset between the nebula semi-major axis and the vector connecting the two quasars in the pair, . To determine and , as described in Arrigoni Battaia et al. (2019a), we first calculated the second order moments of the flux distribution as
@@
with the flux-weighted centroid coordinates and the distance of the point to the centroid, . From this, the Stokes parameters follow as
and were used to calculate the axis ratio, or asymmetry of the nebula,
and the angle of the semi-major axis to the next x- or y-axis,
@@
From , the angle between semi-major axis of the nebula and the connecting line between the quasars can be determined. The determination of positions and as a result most aforementioned values are seeing-limited and this imprecision accounts for the majority of uncertainties. We therefore assumed an error of 1seeing on all distance measurements and obtained errors for and by randomly varying the centroid position 500 times within a circle of radius 1seeing. The minimum and maximum values of the obtained distribution were taken as upper and lower errors. Physical area measurements are also influenced by uncertainties on the redshift. A conservative estimate for this error is 10 percent of the measured area (Arrigoni Battaia et al., 2023b). To avoid taking into account the same nebula twice, we only determined morphologies with respect to the brighter quasar in the pair if both of them are associated with the same nebula or only one quasar is in the FoV (ID 1.1, 4.1, 5.1, 6.1 and 8.1). If there are two clearly distinct nebulae or only one nebula, analysis was performed for each of those individually with respect to the quasar they are associated with (ID 2.1, 2.2, 3.2, 7.1). This analysis was not performed for quasar 7.2, as we require a minimum enclosed area of 2 arcsec2 to be able to sensibly constrain the morphology. Table 3 lists all the aforementioned metrics computed to constrain the morphology of the detected Ly nebulae. On average, the extended emission spans about 90 kpc with a nebula luminosity of 2.8 erg s-1.
@@
We summarize our main findings on nebulae morphologies in Figs. 5 and 6. The first figure shows the axis ratio of each nebula, , against its nebula luminosity , with the dashed red line indicating the median value of . For comparison, we also provide the values reported in Arrigoni Battaia et al. (2019a) for nebulae detected around more luminous single quasars (QSO MUSEUM I) together with their median value (turquoise dash-dotted) and their 25th and 75th percentiles, and the median value of found for nebulae associated with dimmer quasars (dotted gray line, Mackenzie et al. 2021). For a better visualization of the data, the side plots of Fig. 5 provide the distributions of all the points as histograms.
@@
We find that the second least luminous nebula considered for the morphological calculations, ID 2.1, has the most circular shape. The rest of the sample can be divided into two groups: The first one falls well within the distribution of QSO MUSEUM I with moderate asymmetries (), while the second group shows exceptionally high asymmetries (). Spanning almost two orders of magnitude in nebula luminosities and encompassing quasars 2.2, 7.1, and 8, and therefore a multitude of projected separations and magnitudes, there is no clear trend in what drives these asymmetries. Overall, the shape of quasar pair nebulae is more asymmetric than single quasar nebulae with a median value below the 25th percentile of bright quasar nebulae, even though in previous studies, there is a trend of increasing with lower quasar luminosity. In other words, we find that quasars with luminosities as low as those in Mackenzie et al. (2021) but part of a pair are associated with more asymmetric nebulae. The integrated nebula luminosity spans a similar range in both samples, but peaks at a higher value for single quasars, with the brightest nebula having a luminosity of 1.8 erg s-1, and a higher average luminosity of 4.5 erg s-1. Such more asymmetric morphologies in quasar pairs could be due to the presence of intergalactic structures connecting the two quasars in each pair. This tentative evidence of more asymmetric morphologies needs to be confirmed by a larger sample of quasar pair observations.
@@
To investigate this further, Fig. 6 displays the relation between (the maximum distance between the bright quasar in each pair and its associated nebula isophote) and (the angle between the nebula semi-major axis and the line connecting the two quasars). In the case of isolated/unrelated quasars we expect to find a random distribution of angles and no trend with distance. Instead, the two values are correlated with a Spearman correlation coefficient and a chance of coincidental correlation of . We further tested the null hypothesis by randomly orienting the nebulae 50,000 times and calculating the Spearman coefficient for the obtained distributions. In 0.7 % of realizations, the Spearman coefficient is at least as significant as for the data distribution, that is, below -0.7833, and in 0.2 % of realizations, the 4 most extended nebulae all show alignment angles below 20 degrees.
@@
The correlation is not driven by the magnitude of the quasars (color of the dots) or the projected distance between the pair (size of the dots), although at lower projected distances, the offset angle tends to be smaller. Possibly due to the small sample size, the significance of this trend is low, with a -value of 0.5 determined by performing a 2D Kolmogorov–Smirnov test (Peacock, 1983; Fasano & Franceschini, 1987; Press et al., 2007) using the public code ndtest333Written by Zhaozhou Li, https://github.com/syrte/ndtest.
@@ 4.4 Notes on each quasar pair
Quasar 1.1 (Fig. 3) is the brightest object of the sample; it is bright as the quasars targeted in Borisova et al. (2016) or Arrigoni Battaia et al. (2019a). It is also associated with the brightest Ly nebula with a luminosity . Accordingly, its Ly surface brightness profile is an order of magnitude brighter than the median profile of all quasar pairs. Quasar 1.2, on the other hand, is the faintest targeted object and sits at a projected distance of 47.7 kpc from quasar 1.1. The Ly nebula shows a slight elongation between the two quasars, also indicated by the low value of , and the nebular line velocity shift evolves from around quasar 1.1 to close to quasar 1.2. However, this is not reflected in the systemic redshifts of the AGNs (see top-middle panel in Fig. 3), likely due to the associated large uncertainties on these values as only broad lines are available for the redshift determination, leading to typical errors of up to (Shen et al., 2016). The extended Ly emission traces rather quiescent kinematics with an average velocity dispersion of 385 km s-1.
@@
Quasar pair 2, displayed in Fig. 7, is associated with the smallest Ly nebulae in both area and extent. The extended Ly emission around quasar 2.1 is also the only one in the sample that is almost circularly shaped according to the asymmetry parameter , and therefore, the value of (evaluated here to be about 40∘) is not well constrained. The two nebulae do not seem to be physically connected, supported by their redshift, which is close to the systemic redshift of the respective quasar. This might indicate that, although the projected separation is relatively small (100 kpc), the quasars could be at a significant distance from each other. In this framework the pair would be stretched along our line of sight out to a distance of 2.7 Mpc obtained by assuming that the velocity difference between the two quasars ( km s-1) is all due to the Hubble flow.
Alternatively, the two quasars might not be able to illuminate gas in the transverse direction due to obscuration. In other words, their ionization cones do not illuminate the gas in between the two objects. While quasar 2.1 and 2.2. are relatively faint, there have been observations of individual quasars of similar luminosities that show similar or more extended associated nebulae (e.g., compare ID4, ID5, ID6 and ID7 in Mackenzie et al. 2021).
@@
Quasar 3.1 and 3.2 have similar absolute magnitudes and are the second faintest objects (after ID 1.2) observed in this study, but are associated with the third-biggest nebula, which is displayed in Fig. 8. We note that this is true even when taking into account the slightly differing PSF-subtraction method used for this source, which impacts the estimated . Indeed, the standard subtraction of an empirical PSF leads to artifacts close to the quasar, reducing the area enclosed by the 2 isophote, but does not influence different measures of nebula extent like , which are sensitive to large scale emission. The discovered extended emission does not stretch between the quasars, akin to other pairs in the sample, but is concentrated around quasar 3.2 and shows relatively quiescent kinematics ( km s-1) and a shallower surface brightness profile with respect to the median profile of the sample. While the two quasars are similarly luminous, in the spectrum of quasar 3.2, a blueshifted absorption trough is visible in the CIV line (Fig. 1), which may indicate a nuclear outflow (Weymann et al. 1991) possibly helping in illuminating the surrounding gas (Costa et al. 2022).
@@
Displayed in Fig. 10 are the results for quasar 5.1, while quasar 5.2 has not been observed due to the high angular separation. Quasar 5.1 is closely accompanied by two continuum sources first detected in the MUSE data cube and examined in more detail in a companion paper (Herwig et al., 2024). The detected extended emission is found at a significant offset from the quasar, , and is not directly connected to it. The velocity shift in this nebula spans a large range compared to the rest of the sample, with an evolution from to in the east-west direction. The cool gas traced by the Ly emission might be expelled from the interstellar medium of the quasar and companion host galaxies due to tidal interactions.
@@
This object, shown in Figue 13, has already been published in Arrigoni Battaia et al. (2019b) and is reanalyzed in this work to ensure consistency. While the overall shape of the extended emission stays constant in both works, the method employed here is more conservative (pseudo-NB vs channel-by-channel detection) and consequently does not detect low surface brightness emission features. While the nebulae associated with the two quasars are still connected through a gaseous filament, the emission is much less coherent and smaller in area. However, as the emission appears physically connected, we treat it as one nebula in the analysis.